A step index fiber is made with a core of refractive index 1.36, a diameter of 49pm and a fractional refractive index change of 0.025. it is operated at a wavelength of 1.3. Then the numerical aperture of the cable is


1
Expert's answer
2021-04-19T17:26:56-0400

The refractive index is inversely proportional to the wavelength.

Let the refractive index of cladding be n' .

Hence, we know

"\Delta n = \dfrac{n-n'}{n}"


"0.025 = \dfrac{1.36-n'}{1.36}"


"n' = 1.326" = Refractive index of cladding.

We know,

Numerical aperture, "N_a = \sqrt{(n_1^2-n'^2)}"

"n_1 = 1.36"

"n' = 1.326"

"N_a = \sqrt{{1.36^2-1.326^2}}"

"N_a = 0.302"